xx ExternalLibraries/CMake

ExternalLibraries/CMake

Roland Haas Wed Jul 10 15:07:16 CEST 2019

Abstract

Thorn CMakexprovides access to the CMake configuration tool.

1 Using This Thorn

1.1 Interaction With Other Thorns The main product of this thorn is the cmake executable placed in its bin directory.

Since Cactus does not provide a way to ExternalLibraries to modify the search path in $PATH, client thorns need to directly refer to ${CMAKE_DIR}/bin/cmake to use cmake. When making the dependence on this thorn OPTIONAL a construct like

${CMAKE_DIR:+${CMAKE_DIR}/bin/}cmakex$PWD

may be used.

1.2 Examples A typical usage example (for the ADIOS2 thorn) may look like this.

In configuration.ccl have:

OPTIONALxCMake
{
}

In make.code.deps have:

exportxCMAKE_DIR
In build.sh have:

${CMAKE_DIR:+${CMAKE_DIR}/bin/}cmakex-DCMAKE_BUILD_TYPE=${ADIOS2_BUILD_TYPE}x..

2 Parameters

3 Interfaces

General Implements: cmake

4 Schedule This section lists all the variables which are assigned storage by thorn ExternalLibraries/CMake. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE